OpenStudy (anonymous):

Which is a solution to the equation 5x - 2y = -9? A. (2, -4) B. (-1, 2) C. (-2, 5) D. (1, -3)

ganeshie8 (ganeshie8):

where is the second equation ?

OpenStudy (anonymous):

That is all

ganeshie8 (ganeshie8):

oh ok, then simply plugin each point and see which one satisfies

OpenStudy (anonymous):

Okay

ganeshie8 (ganeshie8):

lets see if first option satisfies A. (2, -4) 5x - 2y = -9 plugin x = 2, y = -4 : 4(2) - 2(-4) = -9 8 + 8 = -9 16 = -9 FALSE

ganeshie8 (ganeshie8):

so option A is not ur answer

ganeshie8 (ganeshie8):

similarly, test next option

ganeshie8 (ganeshie8):

B. (-1, 2) 5x - 2y = -9 plugin x = -1, y = 2 5(-1) -2(2) = -9 -5 - 4 = -9 -9 = -9 TRUE

ganeshie8 (ganeshie8):

so, option B is ur answer

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

Okay thanks!!!!